OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (-3, -1) with a slope of 4.

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

To solve this equation, use the POINT - SLOPE formula where we are given a point and a slope. The formula is y−y1=m(x−x1) The slope they give you is 4, so m = 4 The point they give you is (-3,-1) so (x1,y1) x1= -3, y1 = -1 Try not to get confused with the x's and y's. normal x and normal y will always stay the same but x1 and y1 are replaced with the points they give you. Replace the variable with your values and get y - (-1) = (4)(x-(-3)) Then simplify from there, hope that helps!